williams uncle is three times as old as william. in 12 years, williams uncle will be exactly twice his age. how old is william now?
Answer by stanbon(75887)

williams uncle is three times as old as william. in 12 years, williams uncle will be exactly twice his age. how old is william now?
------------
Equations:
U = 3W
----
(U+12)=2(W+12)
---
Substitute for U and solve for W:
(3W+12) = 2W+24
W = 12 yrs. (William's age now)
---------------------
Since U = 3W, U = 36 yrs (the uncle's age now)
====================================================
Cheers,
Stan H.